OpenStudy (anonymous):

write the equation in standard form y^2+4y-8x+4=0

OpenStudy (anonymous):

well y^2+4y+4-8x=0 just rearranged y^2+4y+4=8x y^2+4y+2^2=8x (y+2)^2=8x

OpenStudy (anonymous):

now u tell me ........ this is standard form of equation of wat ?

OpenStudy (anonymous):

well its a parabola .. go on !